OpenStudy (anonymous):

the quotient of n and 2 is less than or equal to 8

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

"the quotient of n and 2" translates to n/2 So "the quotient of n and 2 is less than or equal to 8" means that n/2 <= 8 Now solve for n n/2 <= 8 n <= 8*2 n <= 16

OpenStudy (anonymous):

i should quit for the day

OpenStudy (anonymous):

cause i dont know how

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

can youhelp me graph a solution k/8 less than or equal to negative 5